PLEASE HELP- A TRAIN LEAVES THE STATION AT 6PM GOING WEST AT 80 MI/. ON A PARRALLEL TRACK ANOTHER TRAIN LEAVES 3HRS LATER GOING 100 MI/H. AT WHAT TIME WILL THE 2ND TRAIN CATCH UP WITH THE FIRST TRAIN?
Answer by checkley71(8403)

80t=100(t-3) OR 80t=100t-300 OR 80t-100t=-300 OR -20t=-300 OR t=-300/-20
OR t=-300/-20 OR t=15 hours they will meet.
PROOF: 80*15=100(15-3) or 1200=100*12 or 1200=1200.